OpenStudy (anonymous):

Write an equation for a line in slope-intercept form that passes through the points (3, 4) and (1, -2).

OpenStudy (aroub):

Slope: -2-4/1-3 -6/-2 3 Equation: y=mx+b, m=3 and (3,4) 4=3(3)+b 4=9+b -5=b => y=3x-5